# TSI-Bench: Benchmarking Time Series Imputation

### v1

@@ Abstract
Effective imputation is a crucial preprocessing step for time series analysis. Despite the development of numerous deep learning algorithms for time series imputation, the community lacks standardized and comprehensive benchmark platforms to effectively evaluate imputation performance across different settings. Moreover, although many deep learning forecasting algorithms have demonstrated excellent performance, whether their modeling achievements can be transferred to time series imputation tasks remains unexplored. To bridge these gaps, we develop TSI-Bench, the first (to our knowledge) comprehensive benchmark suite for time series imputation utilizing deep learning techniques. The TSI-Bench pipeline standardizes experimental settings to enable fair evaluation of imputation algorithms and identification of meaningful insights into the influence of domain-appropriate missingness ratios and patterns on model performance. Furthermore, TSI-Bench innovatively provides a systematic paradigm to tailor time series forecasting algorithms for imputation purposes. Our extensive study across 34,804 experiments, 28 algorithms, and 8 datasets with diverse missingness scenarios demonstrates TSI-Bench’s effectiveness in diverse downstream tasks and potential to unlock future directions in time series imputation research and analysis. The source code and experiment logs are available at https://github.com/WenjieDu/Awesome_Imputation.
@@ 1 Introduction
Time series data is widely used in many application domains [1, 2, 3, 4], such as air quality monitoring [5], energy systems [6], traffic [7], and healthcare [8]. However, due to sensor malfunctions, environmental interference, privacy considerations, and other factors, the time series data available for analysis may be incomplete. The absence of complete data can greatly undermine a model’s ability to accurately capture the domain for downstream tasks. By imputing missing values in time series data, the underlying mechanisms leading to missingness can be restored, thereby improving the data completeness and reliability for subsequent analysis and model construction.
@@
The current research landscape in time series imputation is marked by inconsistencies, knowledge gaps, and wide performance disparities reported in literature reviews [9, 10]. These variations arise from differences in data characteristics and dimensionalities, domain-specific missingness patterns, and task-specific requirements. They are further amplified by the heterogeneity of implementation, experiment designs, and evaluation settings. This complicates model and study comparisons, significantly influencing algorithm performance, but remains deeply under-explored [11].
Figure 1: An overview of TSI-Bench.
@@
To bridge existing gaps, we design (to our best knowledge) the first comprehensive time series imputation benchmark, named TSI-Bench. Built on top of an in-house developed imputation suite encompassing 28 highly-performing time-series imputation and forecasting algorithms, TSI-Bench enables efficiently and fairly evaluating newly proposed methods against existing baselines via a standardized ecosystem with a wide range of data preprocessing, missingness simulation and performance evaluation techniques ready for use by researchers to tailor the imputation task according to the needs presented by their datasets, application domains and downstream tasks. Notably, due to forecasting being intertwined with imputation, TSI-Bench also incorporates a systematic paradigm to leverage time series forecasting models for imputation purposes.
The primary contributions of TSI-Bench are as follows:
@@
  • The first comprehensive benchmark for evaluating time series imputation tasks, encompassing 28 ready-to-use algorithms spanning both prediction-based and imputation-based approaches. The benchmark provides eight datasets across four diverse domains (air quality, traffic, electricity, and healthcare) and varying in dimensionality, missingness patterns, and data types.
@@
  • TSI-Bench provides research and application-driven benchmarking perspectives, enabling standardized analysis of the four critical angles of the imputation process we have identified by studying the requirements of various domains: i). the ability to simulate different degrees of missingness and analyze the effect on model performance; ii). the ability to incorporate variations of missingness patterns (e.g., subsequence missing and block missing) in time series according to domain-specific requirements, in addition to single-point missingness in a time series; iii). the types of imputation or forecasting models employed; iv). the expected downstream tasks. Any researcher encountered missing data can use TSI-Bench to explore the suitability of different imputation paradigms to one’s specific problem, dataset, and downstream task.
@@
  • Identification and addressing of limitations inherent in the existing evaluation schemes for TSI to ensure rigorous and equitable comparisons. We build an open-source ecosystem that encompasses a standardized pipeline for data preprocessing, flexible and user-driven missingness simulation, metric utilization, and hyper-parameter tuning. This ecosystem not only facilitates effortless reproduction of our results but also provides users with the tools to assess and integrate their datasets and models into our ecosystem with ease.
@@
Our collective experimental efforts with TSI-Bench assess the efficacy of the different algorithms in a total of 34,804 experiments. Our extensive experiments reveal the following key findings: (1) Different missing patterns and rates significantly influence the performance of imputation methods and none of the 28 algorithms significantly outperforms the others across all settings, emphasizing the importance of the model tuning and data processing capabilities enabled by TSI-Bench ; (2) Forecasting architectures demonstrate effectiveness when used as an imputation backbone; (3) When coupled with domain-informed data processing and simulation of missingness patterns, imputation significantly enhances the data quality and thus boosts downstream task performance.
@@ 2 Preliminaries and Related Work
Problem Definition
A time series sequence of length $L$ can be represented by a matrix $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L})^{\top}\in\mathbf{R}^{D\times L}$, with $D$ being the data dimension. An observation $\mathbf{x}_{t}\in\mathbb{R}^{1\times D}$ is acquired at some timestep $t$. In scenarios involving missing data, each multivariate time series consists of an observed and missing component, i.e., $\mathbf{X}=\{\mathbf{X}^{o},\mathbf{X}^{m}\}$. To encode missingness, we introduce a mask matrix $\mathbf{M}\in\mathbb{R}^{D\times L}$ to represent the existence components of $\mathbf{X}^{o}$, i.e., $\mathbf{M}_{d,l}=1$ if $\mathbf{X}^{o}_{d,l}$ exists, otherwise $\mathbf{M}_{d,l}=0$. We subsequently define the imputation task as follows.
Definition 1 (Time Series Imputation)
Given a reconstructed time series $\bar{X}$ by an imputer, the objective of time series imputation is to obtain an imputed time series,
$\mathbf{\hat{X}}=\mathbf{M}\odot\mathbf{X}^{o}+(1-\mathbf{M})\odot\mathbf{\bar{X}},$ (1)
so as to (i) minimize the discrepancy between the imputed time series $\mathbf{\hat{X}}$ and the real complete time series $\mathbf{X}$ as possible, or (ii) improve the efficacy of downstream tasks when using $\mathbf{\hat{X}}$ compared to using $\mathbf{X}^{o}$ only. $\odot$ denotes the element-wise multiplication.
@@
Deep imputation methods have recently gained popularity in handling missing values in time series data due to their ability to model the nonlinearities and temporal patterns inherent in time series. The literature contains a wide range of deep imputation methods, which can be broadly classified as either predictive or generative.
Predictive imputation methods aim to consistently predict deterministic values for missing components within the time series  [12, 13, 14, 15, 16, 17]. Existing models employ a variety of neural network architectures, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Graph Neural Network (GNN), and attention mechanisms. Highly cited examples include GRU-D [18] and BRITS [19], which achieve improved imputation performance by integrating a time-decay mechanism into RNNs to capture irregularities caused by missing values. TimesNet [13] is a highly-performing CNN model that incorporates a Fast Fourier Transformation to restructure 1D time series into a 2D format. Finally, SAITS [20] is an attention-based model comprising two diagonal-masked self-attention blocks and a weighted-combination block, which leverages attention weights and missingness indicators to enhance imputation precision.
Generative deep imputation models, built upon generative models such as VAEs, GANs, and diffusion models, are distinguished by their ability to generate varied outputs for missing observations, reflecting the inherent uncertainty of the imputation process [21, 22, 23, 24, 25, 26, 27]. Examples include GP-VAE [28], which exploits a Gaussian process as a prior distribution to model temporal dependencies in the incomplete time series; US-GAN [29] increases the complexity of discriminator training by compromising the masking matrix, subsequently leading to the improvement of generator performance through adversarial dynamics; and finally CSDI [30] adapts diffusion models for TSI through a conditioned training strategy by employing a subset of the observed data as conditional inputs to guide the generation process of the missing segments within the time series.
@@
In recent years, numerous deep learning architectures have exhibited exceptional performance in time series forecasting problems. These include Transformer-based structures [31, 32, 33], graph-based structures [34, 35, 36, 37], and MLP-based structures [38, 39, 40].
Transformers excel at time series forecasting due to their ability to model long-range dependencies and capture complex temporal patterns [41]. Examples include Informer [42] and Crossformer [43], which utilize different attention mechanisms to weigh the importance of different time steps dynamically. This is particularly beneficial for time series data, which often contains long-term dependencies and intricate patterns. At the same time, researchers showed that simple linear models can sometimes outperform Transformer architectures in long-term time series forecasting tasks [44]. MLP structures, known for their lower complexity and computational efficiency, have attracted growing interest from researchers. For instance, TSMixer [39] extends the MLP mixer concept from computer vision [45] to time series forecasting by performing mixing operations along both time and feature dimensions [38]. Other innovative MLP-based designs, such as Koopa [46] and FITS [47], further highlight the promising potential of MLP approaches in advancing time series forecasting.
The progress in forecasting models has the potential to be beneficial in imputation problems. Theoretically, forecasting and imputation problems can be regarded as two similar tasks in time series analysis, as both require obtaining temporal representations of existing time series data to predict future values [48] or to impute missing values [10, 49]. Given that time series forecasting is a highly active research area within time series analysis, a framework that enables researchers to directly apply forecasting methods to imputation challenges could potentially lead to the development of more efficient imputation techniques. This, in turn, might further advance research in time series analysis.
@@ 3 The Setup of TSI-Bench
In this section, we introduce the models that we evaluate in the experiments, information of datasets, and relevant implementation details.
@@ 3.1 TSI-Bench Suite
Figure 2: Overview of the designed TSI-Bench suite.
TSI-Bench is a standardized suite that streamlines the imputation process and downstream tasks and is based on our PyPOTS Ecosystem [50]. PyPOTS is a toolkit providing easy standardized access to imputation, classification, clustering, and forecasting algorithms and is openly available on GitHub: https://github.com/WenjieDu/PyPOTS. PyPOTS covers the entire imputation workflow from data loading, data processing, and model building, to data imputation. Our TSI-Bench workflow is shown in Figure 2. First, data is loaded from our time series database, TSDB, which serves as the data warehouse. The loaded data is then transformed into partially-observed time series (POTS) using PyGrinder, which performs flexible and user-driven missingness simulation. Next, the POTS is processed with BenchPOTS, which provides standard data preprocessing pipelines to benchmark machine learning algorithms on POTS. Subsequently, the preprocessed time series is input into an imputation model implemented in PyPOTS to estimate the missing values, resulting in imputed time series. Finally, the imputed data is offered to downstream models to perform additional data analysis tasks, e.g. classification, regression, and forecasting.
@@ 3.2 Datasets
In TSI-Bench, we select 8 diverse and representative datasets from four domains for experiments, across air quality, traffic, electricity, and healthcare. In air quality domain, the BeijingAir dataset includes hourly data on six pollutants and meteorological variables from 12 Beijing sites [51]. The ItalyAir dataset contains 9358 hourly records of chemical sensor responses and pollutant concentrations from March 2004 to February 2005 [52]. In traffic, the PeMS dataset provides hourly road occupancy rates from San Francisco Bay area freeways. The Pedestrian dataset features automated pedestrian counts in Melbourne for 2017 [53]. In electricity, the Electricity dataset records hourly consumption for 370 clients, while the ETT dataset includes power load and oil temperature data, using the ETTh1 subset [32]. For healthcare, the PhysioNet2012 dataset has 12,000 ICU patient records focusing on predicting in-hospital mortality, despite 80% missing values [54], and the PhysioNet2019 dataset includes clinical data from 40,336 ICU patients, using subset A [55].

### v2

@@ Abstract
Effective imputation is a crucial preprocessing step for time series analysis. Despite the development of numerous deep learning algorithms for time series imputation, the community lacks standardized and comprehensive benchmark platforms to effectively evaluate imputation performance across different settings. Moreover, although many deep learning forecasting algorithms have demonstrated excellent performance, whether their modelling achievements can be transferred to time series imputation tasks remains unexplored. To bridge these gaps, we develop TSI-Bench, the first (to our knowledge) comprehensive benchmark suite for time series imputation utilizing deep learning techniques. The TSI-Bench pipeline standardizes experimental settings to enable fair evaluation of imputation algorithms and identification of meaningful insights into the influence of domain-appropriate missing rates and patterns on model performance. Furthermore, TSI-Bench innovatively provides a systematic paradigm to tailor time series forecasting algorithms for imputation purposes. Our extensive study across 34,804 experiments, 28 algorithms, and 8 datasets with diverse missingness scenarios demonstrates TSI-Bench’s effectiveness in diverse downstream tasks and potential to unlock future directions in time series imputation research and analysis. All source code and experiment logs are released at https://github.com/WenjieDu/Awesome_Imputation.
Figure 1: An overview of TSI-Bench.
@@ 1 Introduction
Time series data is widely used in many application domains (Zhang et al., 2024; Jin et al., 2024; Hamilton, 2020; Liu et al., 2023a), such as air quality monitoring (Liang et al., 2023), energy systems (Zhu et al., 2023), traffic (Zeng et al., 2022), and healthcare (Ibrahim et al., 2021). However, due to sensor malfunctions, environmental interference, privacy considerations, and other factors, the time series data available for analysis may be incomplete. The absence of complete data can greatly undermine a model’s ability to accurately capture the domain for downstream tasks. By imputing missing values in time series data, the underlying mechanisms leading to missingness can be restored, thereby improving the data completeness and reliability for subsequent analysis and model construction.
@@
The current research landscape in time series imputation is marked by inconsistencies, knowledge gaps, and wide performance disparities reported in literature reviews (Fang & Wang, 2020; Wang et al., 2024). These variations arise from differences in data characteristics and dimensionalities, domain-specific missing patterns, and task-specific requirements. They are further amplified by the heterogeneity of implementation, experiment designs, and evaluation settings. This complicates model and study comparisons, significantly influencing algorithm performance, but remains deeply under-explored (Qian et al., 2024).
@@
To bridge existing gaps, we designed the first (to the best of our knowledge) comprehensive time series imputation benchmark, named TSI-Bench. TSI-Bench is built on top of an in-house developed imputation suite, which includes 28 high-performing time series imputation and forecasting algorithms. It enables efficient and fair evaluation of newly proposed methods against existing baselines through a standardized ecosystem. This ecosystem provides a comprehensive set of data preprocessing tools, missingness simulation techniques, and performance evaluation methods. Researchers can use these resources to tailor the imputation task to suit the specific needs of their datasets, application domains, and downstream tasks. Notably, because forecasting is intertwined with imputation, TSI-Bench also incorporates a systematic paradigm to leverage time series forecasting models for imputation purposes. A comparison of TSI-Bench with prior works is presented in Table 1.
Table 1: A comparison of TSI-Bench and other well-known time series imputation toolkits GluonTS (Alexandrov et al., 2020), Sktime (Löning et al., 2019), imputeTS (Moritz & Bartz-Beielstein, 2017), mice (Van Buuren & Groothuis-Oudshoorn, 2011), Impyute (Law, 2017), and ImputeBench (Khayati et al., 2020).
TSI-Bench (ours) GluonTS Sktime imputeTS mice Impyute ImputeBench
Naive Approaches?
Machine Learning Approaches?
Hyperparameter Optimization?
Missing Pattern Simulations?
Programming Language Python Python Python R R Python C++
Number of Integrated Datasets $\mathbf{172}$ 0 19 6 20 1 10
Number of Integrated Algorithms $\mathbf{39}$ 6 10 10 1 2 19
The primary contributions of TSI-Bench are as follows.
@@
  • The first comprehensive benchmark for evaluating time series imputation, which includes 28 ready-to-use algorithms spanning both imputation-based and forecasting-based approaches. The benchmark utilizes eight datasets across four diverse domains (air quality, traffic, electricity, and healthcare) with various dimensionalities, missing patterns, and data types.
@@
  • TSI-Bench offers both research- and application-driven benchmarking perspectives, enabling a standardized analysis of four critical dimensions of the imputation process, identified through the study of various domain requirements: (1) the capacity to simulate varying degrees of missingness and assess its impact on model performance; (2) the flexibility to incorporate different missing patterns (e.g., subsequence or block missingness) in time series, tailored to domain-specific needs, in addition to single-point missingness; (3) the selection of imputation or forecasting models; and (4) the alignment with expected downstream tasks. Researchers dealing with missing data can leverage TSI-Bench to evaluate the suitability of different imputation approaches for their specific problem, dataset, and downstream applications.
@@
  • Identifying and addressing limitations inherent in existing TSI evaluation schemes to ensure rigorous and equitable comparisons. We build an open-source ecosystem encompassing a standardized data preprocessing pipeline, flexible and user-driven missingness simulation, metric utilization, and hyperparameter tuning. This ecosystem not only facilitates the effortless reproduction of our results, but also provides users with the tools to assess and integrate their datasets and models into our ecosystem with ease.
@@
Our collective experimental efforts with TSI-Bench assess the efficacy of the different algorithms in a total of 34,804 experiments. Our extensive experiments reveal the following key findings: (1) Different missing patterns and rates significantly influence the performance of imputation methods, and none of the 28 algorithms significantly outperforms the others across all settings, emphasizing the importance of the model tuning and data processing capabilities enabled by TSI-Bench; (2) Forecasting architectures demonstrate effectiveness when used as an imputation backbone; (3) Imputation enhances both flexibility and effectiveness across downstream tasks by enabling the broad application of high-quality imputed time series.
@@ 2 Preliminaries and Related Work
Notation
A time series collection of length $T$ can be represented by a matrix $\mathbf{X}=(\bm{x}_{1},\ldots,\bm{x}_{t},\ldots,\bm{x}_{T})^{\top}\in\mathbf{R}^{T\times D}$, with $D$ being the data dimension. An observation $\bm{x}_{t}\in\mathbb{R}^{1\times D}$ is acquired at time step $t$. To encode the presence of missing values in $\mathbf{X}$, we introduce a mask matrix $\mathbf{M}\in\mathbb{R}^{T\times D}$, i.e., ${M}_{t,d}=1$ if ${X}_{t,d}$ exists, otherwise ${M}_{t,d}=0$. We denote the ground truth by $\mathbf{\tilde{X}}\in R^{T\times D}$. Given a reconstructed time series $\mathbf{\bar{X}}$ by an imputer, the objective of time series imputation is to obtain the imputed time series:
$\mathbf{\hat{X}}=\mathbf{M}\odot\mathbf{X}+(1-\mathbf{M})\odot\mathbf{\bar{X}},$ (1)
so as to minimize the discrepancy between the imputed $\mathbf{\hat{X}}$ and the ground truth $\mathbf{\tilde{X}}$ as possible. $\odot$ denotes the element-wise multiplication.
@@
Deep imputation methods have recently gained popularity in handling missing values in time series data due to their ability to model the nonlinearities and temporal patterns inherent in time series. The literature contains a wide range of deep imputation methods, which can be broadly classified as either predictive or generative. Predictive imputation methods aim to consistently predict deterministic values for missing components within the time series (Yoon et al., 2019; Wu et al., 2023; Cini et al., 2022; Marisca et al., 2022; Ma et al., 2019; Bansal et al., 2021). Existing models employ a variety of neural network architectures, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Graph Neural Network (GNN), and attention mechanisms. Highly cited examples include GRU-D (Che et al., 2018a) and BRITS (Cao et al., 2018), which achieve improved imputation performance by integrating a time-decay mechanism into RNNs to capture irregularities caused by missing values. TimesNet (Wu et al., 2023) is a highly-performing CNN model that incorporates a Fast Fourier Transformation to restructure 1D time series into a 2D format. Finally, SAITS (Du et al., 2023) is an attention-based model comprising two diagonal-masked self-attention blocks and a weighted-combination block, which leverages attention weights and missingness indicators to enhance imputation precision.
Generative deep imputation models, built upon generative models such as VAEs, GANs, and diffusion models, are distinguished by their ability to generate varied outputs for missing observations, reflecting the inherent uncertainty of the imputation process (Mulyadi et al., 2021; Kim et al., 2023; Luo et al., 2018; 2019; Alcaraz & Strodthoff, 2023; Chen et al., 2023b; Yang et al., 2024). Examples include GP-VAE (Fortuin et al., 2020), which exploits a Gaussian process as a prior distribution to model temporal dependencies in the incomplete time series; US-GAN (Miao et al., 2021b) increases the complexity of discriminator training by compromising the masking matrix, subsequently leading to the improvement of generator performance through adversarial dynamics; and finally CSDI (Tashiro et al., 2021) adapts diffusion models for TSI through a conditioned training strategy by employing a subset of the observed data as conditional inputs to guide the generation process of the missing segments within the time series.
@@
In recent years, numerous deep learning architectures have exhibited exceptional performance in time series forecasting problems. Such as Transformer-based structures (Wu et al., 2021; Zhou et al., 2021; Nie et al., 2023), GNN-based structures (Cao et al., 2020; Wu et al., 2020; Liu et al., 2022b; Jin et al., 2023), MLP-based structures (Ekambaram et al., 2023; Chen et al., 2023a; Wang et al., 2023a), and diffusion-based structures (Rasul et al., 2021; Tashiro et al., 2021; Shen & Kwok, 2023; Shen et al., 2024).
Transformers excel at time series forecasting due to their ability to model long-range dependencies and capture complex temporal patterns (Wen et al., 2023). Examples include Informer (Zhou et al., 2021) and Crossformer (Zhang & Yan, 2022), which utilize different attention mechanisms to weigh the importance of different time steps dynamically. This is particularly beneficial for time series data, which often contains long-term dependencies and intricate patterns. At the same time, researchers showed that simple models can sometimes outperform Transformer architectures in long-term time series forecasting tasks (Zeng et al., 2023). MLP structures, known for their lower complexity and computational efficiency, have attracted growing interest from researchers. For instance, TSMixer (Chen et al., 2023a) extends the MLP mixer concept from computer vision (Tolstikhin et al., 2021) to time series forecasting by performing mixing operations along both time and feature dimensions (Ekambaram et al., 2023). Other innovative MLP-based designs, such as Koopa (Liu et al., 2023c) and FITS (Xu et al., 2023), further highlight the promising potential of MLP approaches in advancing time series forecasting.
The progress in forecasting models has the potential to be beneficial in imputation problems. Essentially, forecasting and imputation problems can be regarded as two similar tasks in time series analysis, as both require obtaining temporal representations of existing time series data to predict future values (Wen et al., 2019) or to impute missing values (Wang et al., 2024; 2023b). Given that time series forecasting is a highly active research area within time series analysis, a framework that enables researchers to directly apply forecasting methods to imputation challenges could potentially lead to the development of more efficient imputation techniques. This, in turn, might further advance research in time series analysis.
Figure 2: An overview of the designed TSI-Bench suite.
@@ 3 The Setup of TSI-Bench
In this section, we introduce the overview of the models evaluated in the experiments, along with the dataset descriptions and pertinent implementation details.
@@ 3.1 TSI-Bench Suite
TSI-Bench is a standardized suite that streamlines the imputation process and downstream tasks, built upon the PyPOTS ecosystem (Du, 2023). PyPOTS provides easy, standardized access to a variety of algorithms for imputation, classification, clustering, and forecasting. It supports the entire imputation workflow, from data loading and processing to model building and data imputation. The TSI-Bench workflow, depicted in Figure 2, begins with data being loaded from the time series database, TSDB, which serves as the data warehouse. The loaded data is then transformed into partially observed time series (POTS) using PyGrinder, a flexible, user-driven missingness simulation tool. The POTS is subsequently processed by BenchPOTS, which provides standardized data preprocessing pipelines to benchmark machine learning algorithms on POTS. Following preprocessing, the time series is passed through an imputation model implemented in PyPOTS to estimate the missing values, resulting in the imputed time series. Finally, the imputed data is utilised by downstream models for further analysis, such as classification, regression, and forecasting.
@@ 3.2 Datasets
In TSI-Bench, we conduct experiments using eight diverse and representative datasets from four domains: air quality, traffic, electricity, and healthcare. In the air quality domain, the BeijingAir dataset contains hourly data on six pollutants and meteorological variables from 12 sites in Beijing (Zhang et al., 2017). The ItalyAir dataset includes 9,358 hourly records of chemical sensor responses and pollutant concentrations collected from March 2004 to February 2005 (De Vito et al., 2008). In the traffic domain, the PeMS dataset provides hourly road occupancy rates from the San Francisco Bay Area freeways, while the Pedestrian dataset consists of automated pedestrian counts in Melbourne for the year 2017 (Tang et al., 2020). In the electricity domain, the Electricity dataset records hourly electricity consumption for 370 clients, and the ETT dataset includes power load and oil temperature data, with the ETT_h1 subset used for our experiments (Zhou et al., 2021). For healthcare, the PhysioNet2012 dataset comprises 12,000 Intensive Care Unit (ICU) patient records, focusing on in-hospital mortality prediction in the presence of 80% missing values (Silva et al., 2012). Additionally, the PhysioNet2019 dataset includes clinical data from 40,336 ICU patients, using subset A for analysis (Reyna et al., 2020).